Liquidity ratios explain a company’s…..( text and citation needed) 1 The specific Liquidity ratios: Current and & Quick ratios mean what? ….. (see and cite text) 2 Wal-Mart’s Liquidity ratios: what is the three-year trend? 3 Which trend needs elaboration? … To continue this positive trend Wal-Mart should Or …. To address this negative trend Wal-Mart must

WalMart Inc Balance sheet(s) WalMart Inc Income Statement s) Period Ending ..2015 .2014 ..2013 Supplier Purchases) Cost of Goods Sold Selling and General Administrative Total Other Income/Expenses Net Interest Expense Total Sales 485,651,000 476,294,000 468,651,000 365,086,000 358,069,000 352,297,000 Gross Profit 120,565,000 118,225,000 116,354,000 93,418,000 91,35300088,629,000 Operating Poft27,147,000 26,872,000 27,725,000 Cash And Cash Equivalents 9,135,000 7,281,000 Accounts Receivable 6,78000 6,677 Inventory 45,141,000 44858,000 43,803,000 Other Current Assets 2,224,000 113,000 119,000 Earnings Before Interest And Taxes 2034,000 26,753,000 27,539,000 2,461,000 2,335,000 2,249,000 Income Before Tax 24,573,000 24,418,000 25,290,000 8,105,000 7958,000 16,588,000 16,313,000 17,332,000 Total Current Assets 63,278,000 61,185,000 Fixed Assets 116,655,000 117,907,000 116,681,000 20,497,000 5,987,000 Total Assets 2030600 204,751,000 203,105,000 18,102,000 19,510,000 Income Tax Expense Net Income From Continuing Ops Discontinued Operations 7,985,000 285,000 Current Liabilities Net Income 16,303,000 16,169,000 17,280,000 Accounts Payable 58,583,00057,174,000 12,082,000 59,099,000 12,719,000 14,000,000 shares outstanding USE THESE PRICES Price per share Earnings per share 10.00 $9.00 $ 8.00 Price to Earnings Ratio Current Portion Long Term Debt Other Current Liabilities Total Current Liabilities 65,272,000 69,345,000 43,692,000 44,559,000 8,017,000 41,417,000 for ratio Deferred Long Term Liability 8,805,000 4,543,000 7,613,000 Minority Interest Misc Stocks Options Warrants Retained Earnings Other Stockholder Equity Total Liabilities 122,312 126,439,000 323,000 85,777,000 76,566,00072, 2,462,000 2,362,000 7,168,000 1,394,000 7,746,000 1,505,000 Total Stockholder Equity Total Liabilities+Stockholder Equity 203,706,000 204,751,000 203,105,000
Answer #1

The liquidity ratios of a company can be spelled out under two ratio: current ratio and quick ratio.

The current ratio has been designed to estimate the ability of the company to pay its recent or current liabilities out of liquidity or current assets. Formula: Current assets / Current Liabilities

On the other hand, the quick ratio is the second step on liquidity path where we access the most refined liquidity of the company to pay-off the current liabilities. Formula: Current assets – Inventory / Current Liabilities.

Wal-Mart’s liquidity trend are:

RATIOS

2015

2014

2013

Current ratio

63.28/65.27=0.97

61.19/69.34=0.88

59.94/72.01=0.83

Quick ratio

(63.28-45.14)/65.27=0.28

(61.19-44.86)/69.34=0.24

(59.94-43.8)/72.01=0.22

For both the ratios, the trend shows a rising or improving performance by Wal-Mart, but its overall performance to cover the current liabilities has been very weak. The current assets would not be able to cover the whole payment of the current liabilities.

To address the negative current ratio, Wal-Mart has to keep more amount in the liquid form by taking out of the long-term assets or raising more fund.